What is the angle that has supplement 5 times its complement?
Answer by stanbon(75887)

What is the angle that has supplement 5 times its complement?
-----------
Let the angle be "x".
Complement = 90-x
Supplement = 180-x
-----------
Equation:
supp = 5*comp
180-x = 5(90-x)
180-x = 450-5x
4x = 270
x = 67.5 degrees
=======================
Cheers,
Stan H.